OpenStudy (anonymous):

3x-2y:5x-4y= 6:5. Find x:y=? (Note* I got to the part where 14y:15x, don't know if there is even a step like that).

OpenStudy (anonymous):

Someone please help?

OpenStudy (nali):

just equal to each equaiton to its solution as the following: 3x-2y=6 5x-4y=5 then solve one of them for y and then substitute in the other equation 3x-2y=6 -2y=6-3x devide by -2 y=-3+3/2x then subtitute that in the other equaiton 5x-4y=5 5x-4(-3+3/2x) = 5 5x+12-6x=5 -x+12=5 move the 12 to the other side and then devide by negative one to get the value for x. which is 7 then just plug that in the other equaiton that i changed to the y=-3+3/2x = -3+3/2(7) = 7.5